If you find that (3.00x10^-3) moles of H2 gas have a volume of (1.370x10^-1) L under STP conditions, what is the volume of 1 mol under the same conditions?
Answer #1

at STP conditions,

V1/n1 = V2/n2

(1.370x10^-1) / (3.00x10^-3) = V2 / 1 mol

V2 = 45.6667 L for 1 mol of H2 gas